OpenStudy (anonymous):

What is the answer and step by step process to the linear equation? 3x + 4y = 18 x + 3y = 11 Also: Make a table of values for each linear equation. Include at least three rows of the table. Be sure to label each table. Use at least two points to graph each line. Label each point. Label each line.

OpenStudy (anonymous):

\[3x+4y=18\] \[x+3y=11\] Using the substitution method, we get x as: \[x=11-3y\] Put the value of x in equation 1 and, \[3(11-3y)+4y=18\] \[33-9y+4y=18\] \[33-5y=18\] \[33-18=5y\] \[15=5y\] \[15/5=3=y\] Now put the value of y in second equation to find the real value of x, \[x+3(3)=11\] \[x=11-9\] \[x=2\] So we have values as: \[x=2, y=3\] Now just put a different value for x, in one equation and get the y-value, when you have 2 sets of coordinates for each equation graph it :P

OpenStudy (anonymous):

What would it look like on a graph?